If two angles are supplementary and the larger angle is 10 degrees less than four times the smaller angle, what is the difference of measures of the angles?
Answer by mananth(16946)

let smaller angle be x
larger angle = 4x-10
4x-10+x=180
5x=190
x=38 smaller angle
Larger angle = 4x-10 = 152-10 = 142
..
The difference = 142-38= 104